OpenStudy (hba):

If f(g(x))=4x^2-8x and f(x)=x^2-4 then g(x)= ?

OpenStudy (hba):

@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

wait..... sorry i did a mistake

OpenStudy (hba):

I know.

OpenStudy (anonymous):

A convenient solution is to let y=g(x) and f(y)=f(g(x)) \[y^2 - 4 =4x^2 - 8x\] rearrange for y and voila! you get g(x)

OpenStudy (jiteshmeghwal9):

ya @Shadowys is right :)